# Curriculum Learning Meets Directed Acyclic Graph for Multimodal Emotion Recognition

### 27 Feb 2024 v1

@@ Abstract
Emotion recognition in conversation (ERC) is a crucial task in natural language processing and affective computing. This paper proposes MultiDAG+CL, a novel approach for Multimodal Emotion Recognition in Conversation (ERC) that employs Directed Acyclic Graph (DAG) to integrate textual, acoustic, and visual features within a unified framework. The model is enhanced by Curriculum Learning (CL) to address challenges related to emotional shifts and data imbalance. Curriculum learning facilitates the learning process by gradually presenting training samples in a meaningful order, thereby improving the model’s performance in handling emotional variations and data imbalance. Experimental results on the IEMOCAP and MELD datasets demonstrate that the MultiDAG+CL models outperform baseline models. We release the code for MultiDAG+CL and experiments: https://anonymous.4open.science/r/MultiDAG-CL-3D8F.
@@ 1.   Introduction
Multimodal Emotion Recognition in Conversation (Multimodal ERC) classifies emotions in conversation turns using text, audio, and visual cues. By incorporating multiple modalities, it provides a comprehensive representation of emotional expressions, including tone of voice, facial expressions, and body language, resulting in improved accuracy and robustness in emotion recognition compared to traditional unimodal ERC approaches. The MFN Zadeh et al. (2018) synchronizes multimodal sequences using a multi-view gated memory. ICON Hazarika et al. (2018) provides conversational features from modalities through multi-hop memories. The bc-LSTM Poria et al. (2017) leverages an utterance-level LSTM to capture multimodal features. MMGCN Hu et al. (2021) uses a graph-based fusion module to capture intra- and inter-modality contextual features. CTNet Lian et al. (2021) utilizes a transformer-based structure to model interactions among multimodal features. CORECT Nguyen et al. (2023) leverages relational temporal GNNs with cross-modality interaction support, effectively capturing conversation-level interactions and utterance-level temporal relations.
A Directed Acyclic Graph (DAG) is a directed graph without any directed cycles, comprising vertices and edges, where each edge is directed from one vertex to another, ensuring no closed loops. Building upon this concept, Yu et al. (2019) introduced Directed Acyclic Graph Neural Network (DAG-GNN). Additionally, Shen et al. (2021) presented DAG-ERC, a model combining graph-based and recurrence-based neural architectures to capture information flow in long-distance conversations. However, DAG-ERC’s focus has been primarily on unimodal text data, with limited exploration in other modalities. Curriculum Learning (CL), inspired by human learning, progressively introduces more complex concepts starting from a simple initial state. It establishes a sequence of curricula where the best curriculum with the simplest examples is used to train the classifier in each learning round Bengio et al. (2009); Soviany et al. (2022). Curriculum Learning incorporates two key factors: a difficulty measurer to assess the difficulty level of training examples, and a training scheduler to determine the order of example presentation during training. The difficulty measurer assesses the difficulty level of training examples, while the training scheduler determines the order in which examples are presented to the model during training. For the ERC task, Yang et al. (2022) proposes a hybrid Curriculum Learning framework specifically for the textual modality only.
@@ 3.1.   Datasets and Baselines
We evaluate our approach on the following two ERC datasets: IEMOCAP Busso et al. (2008) and MELD Poria et al. (2019). The detailed statistics of the datasets are reported in Table 2. For the data processing, we use the same split as the work in Hu et al. (2021). We compare our method against several state-of-the-art baselines, including unimodal and multimodal learning approaches. (Due to the space limit, they are brief described in Section 1).
The evaluation metrics used are Accuracy (Acc.) and weighted average F1-score (w-F1).

### 08 Mar 2024 v2

@@ Abstract
Emotion recognition in conversation (ERC) is a crucial task in natural language processing and affective computing. This paper proposes MultiDAG+CL, a novel approach for Multimodal Emotion Recognition in Conversation (ERC) that employs Directed Acyclic Graph (DAG) to integrate textual, acoustic, and visual features within a unified framework. The model is enhanced by Curriculum Learning (CL) to address challenges related to emotional shifts and data imbalance. Curriculum learning facilitates the learning process by gradually presenting training samples in a meaningful order, thereby improving the model’s performance in handling emotional variations and data imbalance. Experimental results on the IEMOCAP and MELD datasets demonstrate that the MultiDAG+CL models outperform baseline models. We release the code for MultiDAG+CL and experiments: https://github.com/vanntc711/MultiDAG-CL.
@@ 1.   Introduction
Multimodal Emotion Recognition in Conversation (Multimodal ERC) classifies emotions in conversation turns using text, audio, and visual cues. By incorporating multiple modalities, it provides a comprehensive representation of emotional expressions, including tone of voice, facial expressions, and body language, resulting in improved accuracy and robustness in emotion recognition compared to traditional unimodal ERC approaches. Several models have been proposed to address the task of multimodal ERC. The MFN Zadeh et al. (2018) synchronizes multimodal sequences using a multi-view gated memory. ICON Hazarika et al. (2018) provides conversational features from modalities through multi-hop memories. The bc-LSTM Poria et al. (2017) leverages an utterance-level LSTM to capture multimodal features. MMGCN Hu et al. (2021) uses a graph-based fusion module to capture intra- and inter-modality contextual features. CTNet Lian et al. (2021) utilizes a transformer-based structure to model interactions among multimodal features. CORECT Nguyen et al. (2023) leverages relational temporal GNNs with cross-modality interaction support, effectively capturing conversation-level interactions and utterance-level temporal relations.
A Directed Acyclic Graph (DAG) is a directed graph without any directed cycles, comprising vertices and edges, where each edge is directed from one vertex to another, ensuring no closed loops. Building upon this concept, Yu et al. (2019) introduced Directed Acyclic Graph Neural Network (DAG-GNN). Additionally, Shen et al. (2021) presented DAG-ERC, a model combining graph-based and recurrence-based neural architectures to capture information flow in long-distance conversations. However, DAG-ERC’s focus has been primarily on unimodal text data, with limited exploration in other modalities. Curriculum Learning (CL), inspired by human learning, progressively introduces more complex concepts starting from a simple initial state. It establishes a sequence of curricula where the best curriculum with the simplest examples is used to train the classifier in each learning round Bengio et al. (2009); Soviany et al. (2022). CL incorporates two key factors: a difficulty measurer to assess the difficulty level of training examples, and a training scheduler to determine the order of example presentation during training. The difficulty measurer assesses the difficulty level of training examples, while the training scheduler determines the order in which examples are presented to the model during training. For the ERC task, Yang et al. (2022) proposes a hybrid CL framework specifically for the textual modality only.
@@ 3.1.   Datasets and Baselines
We evaluate our approach on the following two ERC datasets: IEMOCAP Busso et al. (2008) and MELD Poria et al. (2019). The detailed statistics of the datasets are reported in Table 2. For the data processing, we use the same split as the work in Hu et al. (2021). We compare our method against several state-of-the-art baselines, including unimodal and multimodal learning approaches. (Due to the space limit, they are brief described in Section 1). The evaluation metrics used are Accuracy (Acc.) and weighted average F1-score (w-F1).